Find real numbers x and y such that (3x - 4y) + (6x + 2y)i = 1 - 7i
Answer by venugopalramana(3286)

(3x - 4y) + (6x + 2y)i = 1 - 7i
real part on lhs = real part on rhs
imaginary part on lhs = imaginary part on rhs
3x-4y = 1.......(1)
6x+2y= -7.........(2)
2*(1) gives ...6x-8y=2.......(3)
(2) - (3) gives 2y-(-8y) = -7-2
10y = -9
y= -9/10
substituting in (1) we get 3x-4*(-9/10) =1
3x+18/5 =1
3x=1 - 18/5 = (5-18)/5 = -13/5
x= -13/(5*3) = -13/15